# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v2

@@ Abstract
This study investigates the iterative refinement method applied to the solution of linear discrete inverse problems by considering its application to the Tikhonov problem in mixed precision. Previous works on mixed precision iterative refinement methods for the solution of symmetric positive definite linear systems and least-squares problems have shown regularization to be a key requirement when computing low precision factorizations. For problems that are naturally severely ill-posed, we formulate the iterates of iterative refinement in mixed precision as a filtered solution using the preconditioned Landweber method with a Tikhonov-type preconditioner. Through numerical examples simulating various mixed precision choices, we showcase the filtering properties of the method and the achievement of comparable accuracy compared to results computed in double precision as well as another approximate method.
@@
mixed precision, inverse problems, iterative refinement, Tikhonov regularization, preconditioned iterative methods, Landweber
@@ 1 Introduction
[…]
@@
Given the unexpected necessity of regularization in structure leveraging algorithms for SPD linear systems in [19], we investigate the use of IR to solve the Tikhonov problem in mixed precision providing a natural extension to the aformentioned works. To better understand the regularized solutions of ill-posed problems computed using IR we derive a methodology to formulate the iterates applied to the Tikhonov problem as filtered solutions by writing them as a recursive relationship between the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous iterates. Additionally, we demonstrate in our numerical results that mixed precision IR on the Tikhonov problem gives comparable or superior accuracy against results computed in double precision as well as another benchmark which supports its use in modern applications that natively support mixed precision floating-point arithmetic.
@@ 2.2 Landweber Iteration
For a scaled matrix $A$ whose largest singular value is one, only the first filter factor of the $k^{th}$ iteration, $\phi^{(k)}_{1}$ will be $1$, with subsequent filter factors of the same iteration rapidly decaying to numerical zero. It is only when $k$ grows that early filtering values, i.e. $\phi^{(k)}_{j}$ for small $j$, cluster near $1$. Because of this behavior, it can take many iterations to fully capture the dominant right singular vectors that constitute a meaningful approximate solution for inverse problems. One common way to overcome this difficulty is to employ a preconditioner; we comment more on this in Section 3.
@@ 2.3 Iterative Refinement on the Tikhonov Problem
Variants of fixed and mixed precision IR methods are not new - having been investigated consistently since Wilkinson’s first programmed version in 1948 [32]. A historical treatise of the progression of IR for linear problems may be found in [2, 17]. On a generic linear system, IR computes the $k^{th}$ approximate solution, $x^{(k)}$, in three steps: (i) the $(k-1)^{th}$ residual $r^{(k-1)}$ is computed, (ii) the correction equation $Ah^{(k-1)}=r^{(k-1)}$ is solved, and (iii) the current solution is updated $x^{(k)}=x^{(k-1)}+h^{(k-1)}$. The process may be repeated as necessary until a termination criterion is met.
@@
(10) $\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}\hat{r}^{(k-1)}\\ &=x^{(k-1)}_{IR}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(A^{T}A+\alpha^{2}I\right)^{-1}x_{IR}^{(k-1)}.\end{split}$
@@
1  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x_{IR}^{(0)}\in\mathbb{R}^{n},\,\,\alpha>0$
2  Output: $x_{IR}^{(k+1)}\in\mathbb{R}^{n}$
3  for $k=0,1,2,\dots$ do
@@ 7        $x_{IR}^{(k+1)}=x_{IR}^{(k)}+h^{(k)}$
8  end for
From (10) we may observe that the term $\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}\hat{r}^{(k-1)}$ contains the correction term of iterated Tikhonov (9) whose filter factors at the $k^{th}$ iteration take the form
@@ 3 Landweber with a Tikhonov-type Preconditioner
where $D^{2}=\left(\Sigma^{2}+\alpha^{2}\right)$ and $\Sigma^{2}=\Sigma^{T}\Sigma\in\mathbb{R}^{n\times n}$. Thus, one may consider the preconditioning matrix $M$ of (12) as $M=\left(\Sigma^{2}+\alpha^{2}\right)^{1/2}V^{T}=DV^{T}$: the product of a diagonal matrix with the right singular matrix of $A$.
Herein, we utilize the $M$ notation to help simplify our notation throughout this work and to consider the situation that $M^{T}M$ represents an approximate Tikhonov-type preconditioner to $A^{T}A+\alpha^{2}I$. Possible approximations we consider in Section 5 include that of (1) $M^{T}M$ as a low precision SVD approximation and (2) when the structure of $A$ allows for the approximation by a second matrix $C$ whose structure allows for fast and robust matrix-vector products. The former will be our focus when considering IR on the Tikhonov problem in mixed precision. The latter case can arise naturally, for example, in image deblurring problems; see Section 5.
We clarify that there are various possible preconditioning strategies other than Tikhonov that could be used. It should be understood by our usage of the wording ‘preconditioned Landweber’ in this work that we only consider Tikhonov-type preconditioners. For further discussions and other preconditioner types considered for the solution of linear discrete ill-posed problems see [1, 26, 29].
@@
(15) $x^{(k)}_{PL}=\sum_{j=1}^{n}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}$
@@ 4.1 Reformulated Iterative Refinement on the Tikhonov problem
(20) $\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{k-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(k-1-i)}_{IR}.\end{split}$
@@
$\begin{split}&x^{(n+1)}_{IR}=x^{(n)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax^{(n)}_{IR}\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-A\left[x^{(n)}_{PL}-\Gamma\right]\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n)}_{IR}+\underbrace{\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax_{PL}^{(n)}\right)}_{=h^{(n)}}-\left(M^{T}M\right)^{-1}A^{T}A\Gamma-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n)}_{PL}-\Gamma+h^{(n)}-\left(M^{T}M\right)^{-1}A^{T}A\Gamma-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(n)}_{IR}\\ &=x^{(n+1)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{0}x^{(n)}_{IR}\\ &\hskip 14.22636pt-\alpha^{2}\left(M^{T}M\right)^{-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)\sum_{i=0}^{n-1}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(n-1-i)}_{IR}\\ &=x^{(n+1)}_{PL}-\alpha^{2}\left(M^{T}M\right)^{-1}\sum_{i=0}^{n}\left(I-A^{T}A\left(M^{T}M\right)^{-1}\right)^{i}x^{(n-i)}_{IR}.\end{split}$
@@
(21) $\begin{split}x^{(2)}_{IR}&=x^{(2)}_{PL}-\alpha^{2}(M^{T}M)^{-1}\sum_{i=0}^{1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{i}x^{(1-i)}_{IR}\\ &=x^{(2)}_{PL}-\alpha^{2}(M^{T}M)^{-1}\left(I-A^{T}A(M^{T}M)^{-1}\right)^{0}x^{(1)}_{PL}.\end{split}$
@@
$\begin{split}\phi^{(2)}_{j}&=\psi^{(2)}_{j}-\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{0}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)\\ &={\color[rgb]{0,0,0}\psi^{(2)}_{j}-\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\left(1-\left(\frac{\sigma_{M,j}^{2}+\alpha^{2}-\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{1}\right)}\\ \end{split}$
@@
(IR on Tikhonov problem with filter factors)
@@
$x^{(k)}_{IR}=\sum_{j=1}^{n}\left(\psi^{(k)}_{j}-\left(\frac{\alpha^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)\sum_{i=0}^{k-1}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma_{M,j}^{2}+\alpha^{2}}\right)^{i}\phi_{j}^{(k-1-i)}\right)\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}$
with the $j^{th}$ filter factor per $k^{th}$ step given by
@@
We comment that the filter factors of IR on the Tikhonov problem computed in exact arithmetic are exactly those of Tikhonov regularization. This is stated in Corollary 4.4 below without proof as the argument follows from a straight forward mathematical induction argument.
@@ 4.2 Mixed precision IR on the Tikhonov problem
We now turn to the determination of the filter factors of IR when computing in multiple precisions. We extend the result of Theorem 4.3 to the case when the $k^{th}$ iterate and its corresponding filter factors are to be computed in up to three precisions. When discussing computations in mixed precision we match the convention of recent works, e.g., [13, 17], by using the notation
@@
to represent the three precisions used. Here, Pr3 denotes the highest precision that is used in residual updates, Pr2 denotes the ‘working precision’ used for linear systems solves, and Pr1 denotes the precision used to represent the preconditioner $M$. In Section 5 we define the precisions we consider for our numerical experiments.
@@
where we utilize braces to visually denote the usage of mixed precision. This mixed precision variant of Algorithm 1 is provided in Algorithm 2 for reference. Recall that the $k^{th}$ solution of (19) can be written as a recursive relation involving the $k^{th}$ iterate of preconditoned Landweber and the previous $k-1$ iterates of IR on the Tikhonov problem. In Lemma 4.5 we reformulate (19) to write the $k^{th}$ iterate of the IR problem as a filtered solution whose constituent parts are computed in precisions that mimic their partnered forms of (25) and Algorithm 2.
@@
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,x_{IR}^{(0)}\in\mathbb{R}^{n}$
3  Output: $x_{IR}^{(k+1)}\in\mathbb{R}^{n}$
4  Compute $M^{T}M=\left(A^{T}A+\alpha^{2}I\right);$ $[\text{in }\text{Pr}_{1}]$
5  for $k=0,1,2,\dots$ do
6        $r^{(k)}=b-Ax_{IR}^{(k)};$ $[\text{in }\text{Pr}_{3}]$
7        $s^{(k)}=A^{T}r^{(k)}-\alpha^{2}x_{IR}^{(k)};$ $[\text{in }\text{Pr}_{3}]$
8        Solve $\left(M^{T}M\right)h^{(k)}=s^{(k)};$ $[\text{in }\text{Pr}_{2}]$
9        $x_{IR}^{(k+1)}=x_{IR}^{(k)}+h^{(k)};$ $[\text{in }\text{Pr}_{2}]\,\,$
10  end for
@@ 9        $x_{IR}^{(k+1)}=x_{IR}^{(k)}+h^{(k)};$ $[\text{in }\text{Pr}_{2}]\,\,$
$\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}\left(b-Ax_{IR}^{(k-1)}\right)-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}A^{T}r_{PL}^{(k-1)}+\left(M^{T}M\right)^{-1}\Gamma-\alpha^{2}\left(M^{T}M\right)^{-1}x^{(k-1)}_{IR}\\ &=x^{(k-1)}_{IR}+\left(M^{T}M\right)^{-1}\left[A^{T}r_{PL}^{(k-1)}+\Gamma-\alpha^{2}x^{(k-1)}_{IR}\right]\end{split}$
@@
We comment that during the computation of the $k^{th}$ iterate of (26) that the product $\left(M^{T}M\right)\left[x^{(k)}_{PL}-x^{(k-1)}_{PL}\right]$ is computed in Pr3 but $\left(M^{T}M\right)$ is computed in Pr1. Thus, before presenting the mixed precision generalization of Theorem 4.3, we present in Proposition 2 the variable precision case of Proposition 1. In Algorithm 3 we provide the steps for computing preconditioned Landweber where we note that the preconditioner $\left(M^{T}M\right)$ is an input and the working precision is Pr3 (as per Lemma 4.5). The framework for the computations of the filter factors is summarized in Algorithm 4.
Algorithm 3 Variable precision preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x_{PL}^{(0)}=0\in\mathbb{R}^{n},\,\,\alpha>0$,   $M^{T}M$ (in Pr1)
3  Output: $x_{PL}^{(k+1)}\in\mathbb{R}^{n}$
4  for $k=0,1,2,\dots$ do
5        $s^{(k)}=A^{T}\left(b-Ax_{PL}^{(k)}\right)$$[\text{in }\text{Pr}_{3}]$
6        Solve $(M^{T}M)h^{(k)}=s^{(k)}$ $[\text{in }\text{Pr}_{3}]$
7        $x_{PL}^{(k+1)}=x_{PL}^{(k)}+h^{(k)}$ $[\text{in }\text{Pr}_{3}]\,\,$
8  end for
@@ 7        $x_{PL}^{(k+1)}=x_{PL}^{(k)}+h^{(k)}$ $[\text{in }\text{Pr}_{3}]\,\,$
(Variable precision preconditioned Landweber with filter factors)
@@
(28) $x^{(k)}_{PL}=\sum_{j=1}^{n}\left(\psi_{j}^{(k-1)}+\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k-1}\right)\frac{u^{T}_{A,j}b}{\sigma_{A,j}}v_{M,j}$
@@
(29) $\psi_{j}^{(k)}=\psi_{j}^{(k-1)}+\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\left(1-\frac{\sigma_{A,j}^{2}}{\sigma^{2}_{M,j}+\alpha^{2}}\right)^{k-1}\quad\text{for}\,\,\,j=1,2,\dots,n$
under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
@@
$\begin{split}A^{T}r^{(k-1)}&=\left(M^{T}M\right)\left(x^{(k)}_{PL}-x^{(k-1)}_{PL}\right)\\ &=\left(M^{T}M\right)\left[M^{-1}\left(\hat{x}_{PL}^{k}-\hat{x}_{PL}^{k-1}\right)\right]\\ &=\left(M^{T}M\right)\left[M^{-1}\left(\sum_{i=0}^{k-1}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b-\sum_{i=0}^{k-2}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b\right)\right]\\ &=\left(M^{T}M\right)M^{-1}\left(I-\hat{A}^{T}\hat{A}\right)^{k-1}\hat{A}^{T}b.\end{split}$
@@
Algorithm 4 Filter factors of variable precision preconditioned Landweber
1  Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2  Input: $\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\alpha>0$,

### v3

@@ Abstract
This study investigates the iterative refinement method applied to the solution of linear discrete inverse problems by considering its application to the Tikhonov problem in mixed precision. Previous works on mixed precision iterative refinement methods for the solution of symmetric positive definite linear systems and least-squares problems have shown regularization to be a key requirement when computing low precision factorizations. For problems that are naturally severely ill-posed, we formulate the iterates of iterative refinement in mixed precision as a filtered solution using the preconditioned Landweber method with a Tikhonov-type preconditioner. Through numerical examples simulating various mixed precision choices, we showcase the filtering properties of the method and the achievement of comparable working accuracy of discrete inverse problems (i.e., to within a few decimal places in relative error) compared to results computed in double precision as well as another approximate iterative refinement method which we use as a benchmark.
@@
mixed precision, inverse problems, iterative refinement, Tikhonov regularization, preconditioned iterative methods, Landweber method
@@ 1 Introduction
in computing environments where mixed precision floating-point arithmetic is available. Here, $A\in\mathbb{R}^{m\times n}$, $m\geq n$ is a matrix whose singular values decay without significant gap and cluster at the origin (i.e. the matrix is ill-conditioned). We denote the Euclidean norm by $\|\cdot\|$. These types of problems are commonly referred to as linear discrete ill-posed problems. They can arise through the discretization of Fredholm integral equations of the first kind; see [15, 9], but also arise in massive data streaming problems such as the training of the random feature model in machine learning [27] or limited angle tomography problems including, for example, those from medical imaging [7, 25].
@@
Given the unexpected necessity of regularization in structure leveraging algorithms for SPD linear systems in [19], we investigate the use of IR to solve the Tikhonov problem in mixed precision providing a natural extension to the aformentioned works. To better understand the regularized solutions of ill-posed problems computed using IR we derive a methodology to formulate the iterates applied to the Tikhonov problem as filtered solutions by writing them as a recursive relationship between the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous iterates. Additionally, we demonstrate in our numerical results that mixed precision IR on the Tikhonov problem gives comparable working accuracy (i.e., to within a few decimal places in relative error) against results computed in double precision as well as what we term the Approximated Iterative Refinement method (see Section 5.3), which supports its use in modern applications that natively support mixed precision floating-point arithmetic.
@@ 2.2 Landweber Iteration
For a scaled matrix $A$ whose largest singular value is one, only the first filter factor of the $k^{th}$ iteration, $\phi^{(k)}_{1}$ will be $1$, with subsequent filter factors of the same iteration rapidly decaying to numerical zero. It is only when $k$ grows that early filtering values, i.e., $\phi^{(k)}_{j}$ for small $j$, cluster near $1$. Because of this behavior, it can take many iterations to fully capture the dominant right singular vectors that constitute a meaningful approximate solution for inverse problems. One common way to overcome this difficulty is to employ a preconditioner; we comment more on this in Section 3.
@@ 2.3 Iterative Refinement on the Tikhonov Problem
Variants of fixed and mixed precision IR methods are not new - having been investigated consistently since Wilkinson’s first programmed version around 1948 [32]. A historical treatise of the progression of IR for linear problems may be found in [2, 17]. On a generic linear system, IR computes the $k^{th}$ approximate solution, $x^{(k)}$, in three steps: (i) the $(k-1)^{th}$ residual $r^{(k-1)}$ is computed, (ii) the correction equation $Ah^{(k-1)}=r^{(k-1)}$ is solved, and (iii) the current solution is updated $x^{(k)}=x^{(k-1)}+h^{(k-1)}$. The process may be repeated as necessary until a termination criterion is met.
@@
(10) $\begin{split}x^{(k)}_{IR}&=x^{(k-1)}_{IR}+\left(A^{T}A+\alpha^{2}I\right)^{-1}\hat{r}^{(k-1)}\\ &=x^{(k-1)}_{IR}+\left(A^{T}A+\alpha^{2}I\right)^{-1}A^{T}r^{(k-1)}-\alpha^{2}\left(A^{T}A+\alpha^{2}I\right)^{-1}x_{IR}^{(k-1)}.\end{split}$
@@
1Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x_{IR}^{(0)}\in\mathbb{R}^{n},\,\,\alpha>0$
2Output: $x_{IR}^{(k+1)}\in\mathbb{R}^{n}$
3for $k=0,1,2,\dots$ do
@@ 7   $x_{IR}^{(k+1)}=x_{IR}^{(k)}+h^{(k)}$
8end for
From (10) we may observe that the term $\left(A^{T}A+\alpha^{2}I\right)^{-1}\hat{r}^{(k-1)}$ contains the correction term of iterated Tikhonov (9) whose filter factors at the $k^{th}$ iteration take the form
@@ 3 Landweber with a Tikhonov-type Preconditioner
where $D^{2}=\left(\Sigma^{2}+\alpha^{2}I\right)$ and $\Sigma^{2}=\Sigma^{T}\Sigma\in\mathbb{R}^{n\times n}$. Thus, one may consider the preconditioning matrix $M$ of (12) as $M=\left(\Sigma^{2}+\alpha^{2}I\right)^{1/2}V^{T}=DV^{T}$: the product of a diagonal matrix with the right singular matrix of $A$.
Herein, we utilize the $M$ notation to help simplify our notation throughout this work and to consider the situation that $M^{T}M$ represents an approximate Tikhonov-type preconditioner to $A^{T}A+\alpha^{2}I$. Possible approximations we consider in Section 5 include that of (i) $M^{T}M$ as a low precision SVD approximation and (ii) when the structure of $A$ allows for the approximation by a second matrix $C$ whose structure allows for fast and robust matrix-vector products. The former will be our focus when considering IR on the Tikhonov problem in mixed precision. The latter case can arise naturally, for example, in image deblurring problems; see Section 5.
We clarify that there are various possible preconditioning strategies other than Tikhonov that could be used. It should be understood by our usage of the wording ‘preconditioned Landweber’ in this work that we only consider Tikhonov-type preconditioners. For further discussions and other preconditioner types considered for the solution of linear discrete ill-posed problems see e.g., [1, 26, 29].
@@
(15) $\color[rgb]{0,0,0}x^{(k)}_{PL}=\sum_{j=1}^{n}\psi^{(k)}_{j}\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}$}
@@ 4.1 Reformulated Iterative Refinement on the Tikhonov problem
[…]
@@
[…]
@@
[…]
@@
[…]
@@
(IR on the Tikhonov problem with filter factors)
@@
$x^{(k)}_{IR}=\sum_{j=1}^{n}\phi_{j}^{(k)}\frac{u_{A,j}^{T}b}{\sigma_{A,j}}v_{M,j}$
where the $j^{th}$ filter factor per $k^{th}$ step, $\phi_{j}^{(k)}$, is given by
@@
We comment that the filter factors of IR on the Tikhonov problem computed in exact arithmetic when $M^{T}M=A^{T}A+\alpha^{2}I$ are exactly those of Tikhonov regularization. This is stated in Corollary 4.4 below without proof as the argument follows from a straight forward mathematical induction argument.
@@ 4.2 Mixed precision IR on the Tikhonov problem
We now turn to the determination of the filter factors of IR when computing in multiple precisions. We extend the result of Theorem 4.3 to the case when the $k^{th}$ iterate and its corresponding filter factors are to be computed in up to three precisions. When discussing computations in mixed precision we parallel the convention of recent works, e.g., [13, 17], by using the notation
@@
to represent the three precisions used. Here, Pr3 denotes the highest precision that is used in residual updates, Pr2 denotes the ‘working precision’ used for linear systems solves, and Pr1 denotes the precision used to represent the preconditioner $M$. In Section 5 we define the precisions we consider for our numerical experiments. We comment that a finite precision error analysis could be considered to better understand the error accumulation from use of three different precisions; such a treatment is beyond the scope of this work.
@@
where we utilize braces to visually denote the usage of mixed precision. This mixed precision variant of Algorithm 1 is provided in Algorithm 2 for reference. Recall that the $k^{th}$ solution of (19) can be written as a recursive relation involving the $k^{th}$ iterate of preconditoned Landweber and the previous k1k-1 iterates of IR on the Tikhonov problem. In Lemma 4.5 we reformulate (19) to write the $k^{th}$ iterate of the IR problem as a filtered solution whose constituent parts are computed in precisions that mimic their partnered forms of (25) and Algorithm 2.
@@
1Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2Input: $A\in\mathbb{R}^{m\times n},\,\,x_{IR}^{(0)}\in\mathbb{R}^{n}$
3Output: $x_{IR}^{(k+1)}\in\mathbb{R}^{n}$
[…]
5for $k=0,1,2,\dots$ do
[…]
10end for
@@
[…]
@@
We comment that during the computation of the $k^{th}$ iterate of (26) that the product $\left(M^{T}M\right)\left[x^{(k)}_{PL}-x^{(k-1)}_{PL}\right$] is computed in Pr3 but $\left(M^{T}M\right)$ is computed in Pr1. Thus, before presenting the mixed precision generalization of Theorem 4.3, we present in Proposition 2 the mixed precision version of Proposition 1. In Algorithm 3 we provide the steps for computing preconditioned Landweber where we note that the preconditioner $\left(M^{T}M\right)$ is an input and Pr3 is used uniformly (as per Lemma 4.5). The framework for the computations of the filter factors is summarized in Algorithm 4.
Algorithm 3 Mixed precision preconditioned Landweber
1Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2Input: $A\in\mathbb{R}^{m\times n},\,\,b\in\mathbb{R}^{m},\,\,x_{PL}^{(0)}=0\in\mathbb{R}^{n},\,\,\alpha>0$,   $M^{T}M$ (in Pr1)
3Output: $x_{PL}^{(k+1)}\in\mathbb{R}^{n}$
4for $k=0,1,2,\dots$ do
[…]
8end for
@@
(Mixed precision preconditioned Landweber with filter factors)
@@
(28) $\color[rgb]{0,0,0}x^{(k)}_{PL}=\sum_{j=1}^{n}\psi_{j}^{(k)}\frac{u^{T}_{A,j}b}{\sigma_{A,j}}v_{M,j}$}
@@
(29) $\color[rgb]{0,0,0}\psi_{j}^{(k)}=\underbrace{\psi_{j}^{(k-1)}+\sigma_{A,j}^{2}(\underbrace{\sigma^{2}_{M,j}+\alpha^{2}}_{Pr_{1}})^{-1}\Big(1-\sigma_{A,j}^{2}(\sigma^{2}_{M,j}+\alpha^{2})^{-1}\Big)^{k-1}}_{Pr_{3}}$}
for $j=1,2,\dots,n$ under the assumption that $V_{M}^{T}V_{A}=I=V_{A}^{T}V_{M}$.
@@
$\begin{split}A^{T}r^{(k-1)}&=\left(M^{T}M\right)\left(x^{(k)}_{PL}-x^{(k-1)}_{PL}\right)\\ &=\left(M^{T}M\right)\left[M^{-1}\left(\hat{x}_{PL}^{(k)}-\hat{x}_{PL}^{(k-1)}\right)\right]\\ &=\left(M^{T}M\right)\left[M^{-1}\left(\sum_{i=0}^{k-1}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b-\sum_{i=0}^{k-2}\left(I-\hat{A}^{T}\hat{A}\right)^{i}\hat{A}^{T}b\right)\right]\\ &=\left(M^{T}M\right)M^{-1}\left(I-\hat{A}^{T}\hat{A}\right)^{k-1}\hat{A}^{T}b.\end{split}$
@@
Algorithm 4 Filter factors of mixed precision preconditioned Landweber
1Precision Levels: $\text{Pr}_{1}\leq\text{Pr}_{2}\leq\text{Pr}_{3}$
2Input: $\Psi^{(0)}=0\in\mathbb{R}^{n},\,\,\Sigma^{2}_{A}\in\mathbb{R}^{n\times n},\,\,\alpha>0$,